angelie got a high score in her 60 item her score is so close to a perfect score it is divisible by 2 but it leaves 2 remainders when divided by 4?


1
Expert's answer
2021-09-08T11:42:12-0400

Let "a=" a score of Angelie: "a\leq60." Her score is divisible by 2,then "a=2k, 0\leq k\leq 30."

Her score leaves 2 remainders when divided by 4, then "a=4m+2, 0\leq m\leq14."

Her score is so close to a perfect score, then "m=14, a=58."


Angelie got "58."